OpenStudy (anonymous):

Simplify the following expression. sqrt(-17) sqrt(-51)

OpenStudy (anonymous):

\[\sqrt{-51}=i \sqrt{3}*\sqrt{7}\]

OpenStudy (anonymous):

sqrt(17) not sqrt(7)

OpenStudy (anonymous):

so answer: -17sqrt(3)